OpenStudy (anonymous):

x + y = k x - y = k The solution to the system shown is _____. (0, k) (2, 2) (k, 0) got b

OpenStudy (austinl):

x + y = k x - y = k -------- x=k k+y=5 y=0 x-0=k \(\large{\color{blue}{(k,0)}}\)

OpenStudy (austinl):

5+y=k* Silly typo in the 4th line.